# Picky elements, subnormalisers, and character correspondences

### v1

@@ 1. Introduction
Acknowledgements: I thank Alex Moretó and Noelia Rizo for informing me of their conjectures and for several helpful conversations on the topic. I would also like to thank Mandi Schaeffer Fry for comments on an earlier draft and Frank Himstedt for showing me the preprint [37].
@@ 2. Basic observations
The following characterisation turns out to be very useful and could also be taken as definition of subnormaliser for $p$-elements. It shows that the subnormaliser of a $p$-element $x$ is very closely related to the $p$-local structure “around $x$”:
@@
See also Proposition 6.4 for an analogue for algebraic groups. This shows (see also [27, Thm 2.9]):
@@
Is is tempting to define $G$ has almost normal Sylow $p$-subgroups if ${\operatorname{Sub}}_{G}(x)=G$ for all $p$-elements $x\in G$. This is in some sense at the opposite extreme from TI-Sylow $p$-subgroups, for which any $p$-element $x\neq 1$ lies in exactly one Sylow $p$-subgroup, but the two notions coincide when the Sylow $p$-subgroup is normal. An example for $p=2$ is the group SmallGroup(324,37), of the form $3^{3}.{\mathfrak{A}}_{4}$. In fact, similar examples exist for all primes $p$. Non-solvable examples of almost normal Sylow $p$-subgroups are given by the simple groups ${\operatorname{PSL}}_{3}(7)$, ${\operatorname{PSL}}_{3}(13)$, ${\operatorname{PSU}}_{3}(5)$ and ${\operatorname{PSU}}_{3}(11)$ at $p=3$.
@@ 3.1. Unipotent picky elements
Let $T\leq B$ be a maximal torus contained in a Borel subgroup of $G$. Thus $B$ is the group of $F$-fixed points of an $F$-stable Borel subgroup ${\mathbf{B}}\leq{\mathbf{G}}$ with $F$-stable maximal torus ${\mathbf{T}}$ where $T={\mathbf{T}}^{F}$. Then $B,N:={\mathbf{N}}_{G}({\mathbf{T}})$ form a split BN-pair in $G$ with Weyl group $W:=N/T$. Let $\Phi$ be the root system of this BN-pair, and $\Phi^{+},\Delta\subset\Phi$ be the positive respectively simple roots determined by $B$. Now $U:={\mathbf{O}}_{p}(B)$ is a Sylow $p$-subgroup of $G$ and $B$ is its normaliser. Any element in $U$ can be written uniquely as a product of root elements from the root subgroups $U_{\alpha}$, for $\alpha\in\Phi^{+}$, contained in $U$ (see [24, §23]).
@@
Now assume $F$ acts non-trivially on $W$ and hence on $\Delta$. Let $x\in U$ be picky. Then by Proposition 3.1 it is a product involving root elements from all root subgroups for simple roots $\alpha\in\Delta$. Now by [24, Ex. 23.10], for example, the root subgroup $U_{\alpha}$ consists of elements which are products of root elements of ${\mathbf{G}}$ lying in an $F$-orbit of simple roots of ${\mathbf{G}}$, unless $\alpha$ is the image under the orbit map of a pair of roots of ${\mathbf{G}}$ forming a diagram of type $A_{2},B_{2}$ or $G_{2}$. Thus, if we are not in one of the latter cases, $x$ is a product of root elements of ${\mathbf{G}}$ involving all simple roots and hence is regular.
@@ 3.2. Subnormalisers of unipotent elements
Recall that the standard parabolic subgroups of $G$ are exactly the overgroups of the Borel subgroup $B$, and are in natural bijection with the subsets $\Gamma$ of $\Delta$ [24, Prop. 12.2]; we write $P_{\Gamma}$ for the corresponding parabolic subgroup. Let $x\in U$ be unipotent. Then the lattice of standard parabolic subgroups $P$ of $G$ such that $x\in{\mathbf{O}}_{p}(P)$ has a unique maximal element with respect to inclusion, which we denote $P(x)$. Indeed, writing $x$ as a product of root elements $u_{\alpha}\in U_{\alpha}$, with $\alpha\in\Phi^{+}$, we have $P(x)=P_{\Gamma}$ for $\Gamma:=\{\alpha\in\Delta\mid u_{\alpha}=1\}$. The Chevalley commutator formula then shows that $P(x)=P(x^{u})$ for all $u\in U$.
@@
Since $x$ is subnormal in $P(x)$ by definition, it suffices to see that any $H\leq G$ such that $\langle x\rangle{\,\triangleleft\triangleleft\,}H$ lies in $P(x)$. Let $H$ be such a subgroup. In particular $x\in{\mathbf{O}}_{p}(H)$ by Lemma 2.5. By the Borel–Tits theorem [24, Thm 26.5] there is a parabolic subgroup $H^{\prime}\geq H$ such that $x\in{\mathbf{O}}_{p}(H^{\prime})$, so we may assume $H$ is parabolic. Furthermore, we may replace $H$ by any overgroup $H^{\prime}$ with $x\in{\mathbf{O}}_{p}(H^{\prime})$. As any such overgroup is again parabolic, we may assume $H$ is parabolic and maximal with respect to $x\in{\mathbf{O}}_{p}(H)$.
@@
Let $\alpha\in\Delta_{1}$ be a simple root and $P_{\alpha}:=P_{\{\alpha\}}$ the corresponding standard parabolic subgroup of $G$. By assumption $x$ lies in the unipotent radical of $P_{\alpha}$, so $\langle x\rangle$ is subnormal in $P_{\alpha}$. The claim follows since $\langle P_{\alpha}\mid\alpha\in\Delta_{1}\rangle=P_{\Delta_{1}}$ (see [24, Prop. 12.2]). The final assertion is now obvious. ∎
@@
For regular elements this is Theorem 3.2. For non-regular elements, we first consider $G={\operatorname{SL}}_{n}(q)$. For $n=2$ the only non-regular unipotent element is the identity, for $n=3$ the non-trivial non-regular unipotent elements have Jordan type $(2,1)$ and thus are conjugate to a root element for the non-simple positive root, so ${\operatorname{Sub}}_{G}(x)=G$ by Proposition 3.5. Now assume $n\geq 4$. Let $x\in G$ be a non-regular unipotent element. Then $x$ has at least two Jordan blocks, so up to conjugation, $x$ can be written as a product of simple root elements in which at least one simple root $\alpha$ is missing. If this $\alpha$ lies at an end of the Dynkin diagram, then $x^{s_{\alpha}}$ is a conjugate in $B$ in which the simple root next to the end node is missing. So we may assume $\alpha\neq\alpha_{1},\alpha_{n-1}$. Let $P_{1},P_{2}$ be the end-node standard parabolic subgroups of $G$, corresponding to $\Delta\setminus\{\alpha_{1}\}$, $\Delta\setminus\{\alpha_{n-1}\}$ respectively. Let $L_{i}$ be the corresponding standard Levi subgroups, so $L_{i}\cong P_{i}/U_{i}$, with $U_{i}={\mathbf{O}}_{p}(P_{i})$ the unipotent radical of $P_{i}$. Then for $i\in\{1,2\}$, the image $\bar{x}$ of $x$ in $L_{i}$ is a product of root elements not involving the simple root $\alpha$ of $L_{i}$, hence not regular, so by induction ${\operatorname{Sub}}_{L_{i}}(\bar{x})=L_{i}$. (Note that $L_{i}\cong{\operatorname{GL}}_{n-1}(q)$, containing ${\operatorname{SL}}_{n-1}(q)$ as a normal subgroup of $p^{\prime}$-index.) As $U_{i}$ is a $p$-group, ${\operatorname{Sub}}_{P_{i}}(x)=P_{i}$ for $i=1,2$ by Lemma 2.11. Since $\langle P_{1},P_{2}\rangle=G$ by [24, Prop. 12.2] this achieves the proof.
@@ 3.3. Non-simply laced and twisted types
From now on, we number the simple roots in $\Delta$, for ${\mathbf{G}}$ a simple group, as in [24, Tab. 9.1, Tab. 23.2], and we write $P_{i}:=P_{\alpha_{i}}:=P_{\{\alpha_{i}\}}$ for $\alpha_{i}\in\Delta$.
@@
Let us start with $G=B_{2}(q)$. Here representatives for the unipotent conjugacy classes are given in [9, 31]. With the criterion in Proposition 3.5 we see that for $p=2$ there cannot be a class (apart from the regular ones) with ${\operatorname{Sub}}_{G}(x)<G$, and for $p>2$ only the class with representative denoted $A_{22}$ in [31] could possibly have that property. Representatives for the unipotent conjugacy classes of $G_{2}(q)$ are given in [6, 10, 11], and again by Proposition 3.5 it follows that only the class of $G_{2}(2^{f})$ denoted $A_{4}$ in [11] and the class of $G_{2}(p^{f})$, $p\geq 5$, with representative $u_{4}$ in [6] could have ${\operatorname{Sub}}_{G}(x)<G$. For $G={}^{3}\!D_{4}(q)$, consulting the tables of class representatives in [12, 16], we see that we only need to consider the classes with representatives $u_{\beta}(1)u_{2\alpha+\beta}(a)$ and $u_{\alpha}(1)u_{\alpha+\beta}(1)$. Now note that conjugating $u_{\beta}(1)u_{2\alpha+\beta}(a)$ with the simple reflection $s_{\alpha}$ gives an element not involving any simple root element. So we are left with the class $D_{4}(a_{1})$ with representative $u_{\alpha}(1)u_{\alpha+\beta}(1)$. Finally, for $G={}^{2}\!F_{4}(q^{2})$ by [30, Tab. II] apart from the picky cases identified in Theorem 3.2 only the representatives $u_{9},u_{10},u_{11},u_{12}$ involve one of the simple root elements.
@@
Let $P$ be the standard maximal parabolic subgroup of $G$ of type $A_{n-1}$, that is, the stabiliser of a maximal isotropic subspace in the natural representation, and $Q$ the standard maximal parabolic subgroup of type $B_{n-1}$. We claim that if (up to conjugation) ${\operatorname{Sub}}_{P}(x)=P$ then ${\operatorname{Sub}}_{G}(x)=G$. Indeed, assume ${\operatorname{Sub}}_{P}(x)=P$ and let $\bar{x}$ be the image of $x$ in the standard Levi subgroup $L\cong{\operatorname{GL}}_{n}(q)$ of $P$. By Proposition 3.6, $\bar{x}$ is not regular, so (up to conjugation) its image $y$ in the Levi factor of a standard parabolic subgroup $L_{1}$ of $L$ of type ${\operatorname{GL}}_{n-1}(q)$ is not regular either. Now, $L_{1}$ is a standard Levi subgroup of the standard Levi subgroup $M$ of $Q$ of type ${\operatorname{SO}}_{2n-1}(q)$, and hence by induction ${\operatorname{Sub}}_{M}(y)=M$. (The induction base is given by the case $52n-1=$ 5 from Proposition 3.7.) But then ${\operatorname{Sub}}_{Q}(x)=Q$, and since $P$ is maximal in $G$ and $Q$ is distinct from $P$, we have ${\operatorname{Sub}}_{G}(x)=G$.
@@
Let $G=F_{4}(q)$ and $x\in G$ be unipotent. Then ${\operatorname{Sub}}_{G}(x)=G$ unless $x$ is regular, or $q$ is odd and $x$ is in class $F_{4}(a_{1})$, with $|{\mathbf{C}}_{G}(x)|=2q^{6}$ and representative $x_{24}$ in [32, Tab. 5 resp.  6], where ${\operatorname{Sub}}_{G}(x_{24})=P_{\{3,4\}}$.
@@
Representatives for the unipotent conjugacy classes of $G$ as well as for a Levi subgroup $L$ of $G$ of type $B_{3}$ were determined in [29, 32]. First assume $p=2$. We claim that only the regular unipotent elements in $L$ have subnormaliser strictly smaller than $L$. Indeed, seven of the class representatives given in [29, Prop. 2.1] do not involve any simple root element, and the other three non-regular ones can be conjugated by a simple reflection to elements $x$ which only involve the 3rd simple root of $L$. Then $x$ has trivial image in the standard Levi subgroup of $L$ of type $B_{2}$, and image involving just one simple root in the standard Levi subgroup of type $A_{2}$, so ${\operatorname{Sub}}_{L}(x)=L$ by Proposition 3.5.
Now we discuss the unipotent class representatives of $G$ from [29, Thm 2.1]. The ones labelled $x_{0}$ through $x_{19}$ do not involve any simple root element, and similarly for $x_{27},x_{28}$. For elements $x\in\{x_{22},\ldots,x_{26}\}$ only one of the short simple roots occurs, so the images of $x$ in standard Levi subgroups of type $A_{2}$ and $B_{3}$ are non-regular and we conclude by Proposition 3.5 and the previous paragraph. Similarly for $x=x_{20},x_{21}$ the image in a Levi of type $C_{3}$ is trivial, and in a Levi of type $B_{3}$ is non-regular. By Lemma 3.8 this leaves us with the elements $x_{29}$ and $x_{30}$, of type $F_{4}(a_{1})$. The given representatives show that these elements are regular in the subsystem subgroup of type $B_{4}$. Conjugating $x=x_{29}$ or $x_{30}$ with the simple reflection not belonging to the $B_{3}$-subsystem, we obtain an element which only involves simple root elements for the 2nd and 3rd simple root. By the first paragraph, the subnormaliser of that element contains the standard parabolic subgroup of type $B_{3}$, but by Proposition 3.6 also the one of type $A_{2}$, hence it is all of $G$.
If $p$ is odd, Proposition 3.7 shows that apart from the regular classes, only the unipotent class of $L$ with representative $z_{8}$ (in the notation of [32, Tab. 3]) has proper subnormaliser. From [32, Tab. 5 and 6] we conclude that again at most the classes of $G$ of type $F_{4}(a_{1})$ with representatives $x_{23},x_{24}$ might have a proper subnormaliser. Conjugating $x=x_{23},x_{24}$ by the simple reflection not in the $B_{3}$-subsystem we obtain elements whose image in $L$ equals $z_{7},z_{8}$ respectively. As observed above, ${\operatorname{Sub}}_{L}(z_{7})=L$ and so ${\operatorname{Sub}}_{G}(x_{23})=G$ by the usual argument. On the other hand, we have ${\operatorname{Sub}}_{L}(z_{8})$ is the standard parabolic subgroup corresponding to the 3rd node of the diagram of $B_{3}$, which is also the 3rd node of the diagram of $F_{4}$. Thus the set of standard parabolic subgroups of $G$ containing $x_{24}$ in their radical has the unique maximal element $P_{\{3,4\}}$, and so ${\operatorname{Sub}}_{G}(x_{24})=P_{\{3,4\}}$ by Lemma 2.10. ∎
@@
Let $G={\operatorname{SU}}_{2m}(q)$ and $x\in G$ be unipotent. Then ${\operatorname{Sub}}_{G}(x)=G$ unless $x$ is regular, or $x$ has Jordan type $(2m-1,1)$, when ${\operatorname{Sub}}_{G}(x)\sim_{G}P_{m}$.

### v2

@@ 1. Introduction
Acknowledgements: I thank Alex Moretó and Noelia Rizo for informing me of their conjectures and for several enlightening conversations on the topic. I would also like to thank Frank Himstedt for showing me the preprint [40], Martin van Beek for providing the example in Remark 2.11, Frank Lübeck for computing the elements of order 9 in $E_{7}(2)$ and $E_{8}(2)$, and Mandi Schaeffer Fry for comments on an earlier draft.
@@ 2. Basic observations
The following characterisation turns out to be very useful and could also be taken as definition of subnormaliser for $p$-elements. It shows that the subnormaliser of a $p$-element $x$ is very closely related to the $p$-local structure “around xx”:
@@
See also Proposition 6.4 for an analogue for algebraic groups. This shows (see also [30, Thm 2.9]):
@@
The subnormaliser of a $p$-element controls its fusion in a Sylow subgroup; Moretó and Rizo had obtained a different proof of this fact based upon [5, Prop. 2.1].
Lemma 2.9.
Let $P$ be a Sylow $p$-subgroup of $G$ and $x\in P$. Then ${\operatorname{Sub}}_{G}(x)$ is generated by the elements $g\in G$ with $x^{g}\in P$.
Proof.
Let $x^{g}\in P$. Then $x\in{}^{g}\!P$ whence $^{g}\!P\leq{\operatorname{Sub}}_{G}(x)$. Since all Sylow $p$-subgroups of ${\operatorname{Sub}}_{G}(x)$ are conjugate, there is $h\in{\operatorname{Sub}}_{G}(x)$ with $^{g}\!P=P^{h}$ and so $hg\in{\mathbf{N}}_{G}(P)\leq{\operatorname{Sub}}_{G}(x)$, giving $g\in{\operatorname{Sub}}_{G}(x)$.
Conversely, let $R:=\langle g\in G\mid x^{g}\in P\rangle$. Clearly, ${\mathbf{N}}_{G}(P)\leq R$. Assume $x\in{}^{h}\!P$ for some $h\in G$. Then $x,x^{h}\in P$ and thus $h\in R$. Thus, $R$ contains the normalisers of all Sylow $p$-subgroups of $G$ containing $x$ and hence ${\operatorname{Sub}}_{G}(x)$ by Proposition 2.6. ∎
Corollary 2.10.
Let $x\in G$ be a $p$-element. Then $x$ is $G$-conjugate to $y\in{\operatorname{Sub}}_{G}(x)$ if and only if $x,y$ are already ${\operatorname{Sub}}_{G}(x)$-conjugate. In fact, ${\operatorname{Sub}}_{G}(x)$ is the smallest subgroup of $G$ containing both ${\mathbf{C}}_{G}(x)$ and a Sylow $p$-subgroup of $G$ with this property.
Proof.
If $x,y\in{\operatorname{Sub}}_{G}(x)$ are $G$-conjugate, then up to conjugation in ${\operatorname{Sub}}_{G}(x)$ we may assume they lie in a common Sylow $p$-subgroup $P$ (of ${\operatorname{Sub}}_{G}(x)$). Then the first claim follows by Lemma 2.9. For the second, let $R\leq G$ be a subgroup containing ${\mathbf{C}}_{G}(x)$ and a Sylow $p$-subgroup $P$ of $G$ (which we may assume contains $x$) with the stated property. Let $g\in G$ with $x^{g}\in P$ and thus $x,x^{g}$ are $R$-conjugate by assumption. That is, there exists $r\in R$ with $x=x^{gr}$, whence $gr\in{\mathbf{C}}_{G}(x)\leq R$ and so $g\in R$. Lemma 2.9 then shows ${\operatorname{Sub}}_{G}(x)\leq R$. ∎
Remark 2.11.
It is tempting to ask whether the $p$-fusion system of the subnormaliser of a $p$-element of $G$ is determined by the $p$-fusion system of $G$. But this is not the case. I’m indebted to Martin van Beek for the following counter-example: the 3-fusion categories of $M_{12}$ and ${\operatorname{PSL}}_{3}(3)$ are known to be isomorphic. Yet, as we will show, ${\operatorname{PSL}}_{3}(3)$ has a (regular unipotent) picky 3-element, while the corresponding 3-elements in $M_{12}$ have centraliser isomorphic to $3\times{\mathfrak{A}}_{4}$ and subnormaliser $M_{12}$.
Proposition 2.12.
Let $P\leq G$ be a Sylow $p$-subgroup and $x\in P$. If $P$ is abelian then ${\operatorname{Sub}}_{G}(x)=\langle{\mathbf{C}}_{G}(x),{\mathbf{N}}_{G}(P)\rangle$.
Proof.
By Burnside’s theorem, in this case ${\mathbf{N}}_{G}(P)$ controls fusion of elements in $P$, so the claim follows from Corollary 2.10.
Alternatively, let $H:=\langle{\mathbf{C}}_{G}(x),{\mathbf{N}}_{G}(P)\rangle$. Since $x$ is subnormal in both ${\mathbf{C}}_{G}(x)$ and ${\mathbf{N}}_{G}(P)$ we have $H\leq{\operatorname{Sub}}_{G}(x)$. On the other hand, as $P$ is abelian, ${\mathbf{C}}_{G}(x)$, and hence $H$, contains all Sylow $p$-subgroups of $G$ containing $x$. As $H$ contains ${\mathbf{N}}_{G}(P)$, and all Sylow $p$-subgroups of $H$ are $H$-conjugate, it even contains the normalisers of all Sylow $p$-subgroups of $G$ containing $x$. Hence ${\operatorname{Sub}}_{x}(G)\leq H$ by Proposition 2.6, and we are done. ∎
Observe that Proposition 2.6, Lemma 2.9, Corollary 2.10 and Proposition 2.12 hold more generally with $x$ replaced by any $p$-subgroup $H$ of $G$ and ${\mathbf{C}}_{G}(x)$ by ${\mathbf{N}}_{G}(H)$, with identical proofs.
@@
Is is tempting to define $G$ has almost normal Sylow $p$-subgroups if ${\operatorname{Sub}}_{G}(x)=G$ for all $p$-elements $x\in G$. This is in some sense at the opposite extreme from TI-Sylow $p$-subgroups, for which any $p$-element $x\neq 1$ lies in exactly one Sylow $p$-subgroup, but the two notions coincide when the Sylow $p$-subgroup is normal. An example with non-normal Sylow $p$-subgroup for $p=2$ is the group SmallGroup(324,37), of the form $3^{3}.{\mathfrak{A}}_{4}$. In fact, similar examples exist for all primes $p$. Non-solvable examples of almost normal Sylow $p$-subgroups are given by the simple groups ${\operatorname{PSL}}_{3}(7)$, ${\operatorname{PSL}}_{3}(13)$, ${\operatorname{PSU}}_{3}(5)$ and ${\operatorname{PSU}}_{3}(11)$ at $p=3$.
@@ 3.1. Unipotent picky elements
Let $T\leq B$ be a maximal torus contained in a Borel subgroup of $G$. Thus $B$ is the group of $F$-fixed points of an $F$-stable Borel subgroup ${\mathbf{B}}\leq{\mathbf{G}}$ with $F$-stable maximal torus $\mathbf{T}$} where $T={\mathbf{T}}^{F}$. Then $B,N:={\mathbf{N}}_{G}({\mathbf{T}})$ form a split BN-pair in $G$ with Weyl group $W:=N/T$. Let $\Phi$ be the root system of this BN-pair, and $\Phi^{+},\Delta\subset\Phi$ be the positive respectively simple roots determined by $B$. Now $U:={\mathbf{O}}_{p}(B)$ is a Sylow $p$-subgroup of $G$ and $B$ is its normaliser. Any element in $U$ can be written uniquely as a product of root elements from the root subgroups $U_{\alpha}$, for $\alpha\in\Phi^{+}$, contained in $U$ (see [27, §23]).
@@
Now assume $F$ acts non-trivially on $W$ and hence on $\Delta$. Let $x\in U$ be picky. Then by Proposition 3.1 it is a product involving root elements from all root subgroups for simple roots $\alpha\in\Delta$. Now by [27, Ex. 23.10], for example, the root subgroup $U_{\alpha}$ consists of elements which are products of root elements of $\mathbf{G}$} lying in an $F$-orbit of simple roots of $\mathbf{G}$}, unless $\alpha$ is the image under the orbit map of a pair of roots of $\mathbf{G}$} forming a diagram of type $A_{2},B_{2}$ or $G_{2}$. Thus, if we are not in one of the latter cases, $x$ is a product of root elements of $\mathbf{G}$} involving all simple roots and hence is regular.
@@ 3.2. Subnormalisers of unipotent elements
Recall that the standard parabolic subgroups of $G$ are exactly the overgroups of the Borel subgroup $B$, and are in natural bijection with the subsets $\Gamma$ of $\Delta$ [27, Prop. 12.2]; we write $P_{\Gamma}$ for the corresponding parabolic subgroup. Let $x\in U$ be unipotent. Then the lattice of standard parabolic subgroups $P$ of $G$ such that $x\in{\mathbf{O}}_{p}(P)$ has a unique maximal element with respect to inclusion, which we denote $P(x)$. Indeed, writing $x$ as a product of root elements $u_{\alpha}\in U_{\alpha}$, with $\alpha\in\Phi^{+}$, we have $P(x)=P_{\Gamma}$ for $\Gamma:=\{\alpha\in\Delta\mid u_{\alpha}=1\}$. The Chevalley commutator formula then shows that $P(x)=P(x^{u})$ for all $u\in U$.
@@
Since $x$ is subnormal in $P(x)$ by definition, it suffices to see that any $H\leq G$ such that $\langle x\rangle{\,\triangleleft\triangleleft\,}H$ lies in $P(x)$. Let $H$ be such a subgroup. In particular $x\in{\mathbf{O}}_{p}(H)$ by Lemma 2.5. By the Borel–Tits theorem [27, Thm 26.5] there is a parabolic subgroup $H^{\prime}\geq H$ such that $x\in{\mathbf{O}}_{p}(H^{\prime})$, so we may assume $H$ is parabolic. Furthermore, we may replace $H$ by any overgroup $H^{\prime}$ with $x\in{\mathbf{O}}_{p}(H^{\prime})$. As any such overgroup is again parabolic, we may assume $H$ is parabolic and maximal with respect to $x\in{\mathbf{O}}_{p}(H)$.
@@
Let $\alpha\in\Delta_{1}$ be a simple root and $P_{\alpha}:=P_{\{\alpha\}}$ the corresponding standard parabolic subgroup of $G$. By assumption $x$ lies in the unipotent radical of $P_{\alpha}$, so $\langle x\rangle$ is subnormal in $P_{\alpha}$. The claim follows since $\langle P_{\alpha}\mid\alpha\in\Delta_{1}\rangle=P_{\Delta_{1}}$ (see [27, Prop. 12.2]). The final assertion is now obvious. ∎
@@
For regular elements this is Theorem 3.2. For non-regular elements, we first consider $G={\operatorname{SL}}_{n}(q)$. For $n=2$ the only non-regular unipotent element is the identity, for $n=3$ the non-trivial non-regular unipotent elements have Jordan type $(2,1)$ and thus are conjugate to a root element for the non-simple positive root, so ${\operatorname{Sub}}_{G}(x)=G$ by Proposition 3.5. Now assume $n\geq 4$. Let $x\in G$ be a non-regular unipotent element. Then $x$ has at least two Jordan blocks, so up to conjugation, $x$ can be written as a product of simple root elements in which at least one simple root $\alpha$ is missing. If this $\alpha$ lies at an end of the Dynkin diagram, then $x^{s_{\alpha}}$ is a conjugate in $B$ in which the simple root next to the end node is missing. So we may assume $\alpha\neq\alpha_{1},\alpha_{n-1}$. Let $P_{1},P_{2}$ be the end-node standard parabolic subgroups of $G$, corresponding to $\Delta\setminus\{\alpha_{1}\}$, $\Delta\setminus\{\alpha_{n-1}\}$ respectively. Let $L_{i}$ be the corresponding standard Levi subgroups, so $L_{i}\cong P_{i}/U_{i}$, with $U_{i}={\mathbf{O}}_{p}(P_{i})$ the unipotent radical of $P_{i}$. Then for $i\in\{1,2\}$, the image $\bar{x}$ of $x$ in $L_{i}$ is a product of root elements not involving the simple root $\alpha$ of $L_{i}$, hence not regular, so by induction ${\operatorname{Sub}}_{L_{i}}(\bar{x})=L_{i}$. (Note that $L_{i}\cong{\operatorname{GL}}_{n-1}(q)$, containing ${\operatorname{SL}}_{n-1}(q)$ as a normal subgroup of $p^{\prime}$-index.) As $U_{i}$ is a $p$-group, ${\operatorname{Sub}}_{P_{i}}(x)=P_{i}$ for $i=1,2$ by Lemma 2.15. Since $\langle P_{1},P_{2}\rangle=G$ by [27, Prop. 12.2] this achieves the proof.
@@ 3.3. Non-simply laced and twisted types
From now on, we number the simple roots in $\Delta$, for $\mathbf{G}$} a simple group, as in [27, Tab. 9.1, Tab. 23.2], and we write $P_{i}:=P_{\alpha_{i}}:=P_{\{\alpha_{i}\}}$ for $\alpha_{i}\in\Delta$.
@@
Let us start with $G=B_{2}(q)$. Here representatives for the unipotent conjugacy classes are given in [9, 34]. With the criterion in Proposition 3.5 we see that for $p=2$ there cannot be a class (apart from the regular ones) with ${\operatorname{Sub}}_{G}(x)<G$, and for $p>2$ only the class with representative denoted $A_{22}$ in [34] could possibly have that property. Representatives for the unipotent conjugacy classes of $G_{2}(q)$ are given in [6, 10, 11], and again by Proposition 3.5 it follows that only the class of $G_{2}(2^{f})$ denoted $A_{4}$ in [11] and the class of $G_{2}(p^{f})$, $p\geq 5$, with representative $u_{4}$ in [6] could have ${\operatorname{Sub}}_{G}(x)<G$. For $G={}^{3}\!D_{4}(q)$, consulting the tables of class representatives in [12, 16], we see that we only need to consider the classes with representatives $u_{\beta}(1)u_{2\alpha+\beta}(a)$ and $u_{\alpha}(1)u_{\alpha+\beta}(1)$. Now note that conjugating $u_{\beta}(1)u_{2\alpha+\beta}(a)$ with the simple reflection $s_{\alpha}$ gives an element not involving any simple root element. So we are left with the class $D_{4}(a_{1})$ with representative $u_{\alpha}(1)u_{\alpha+\beta}(1)$. Finally, for $G={}^{2}\!F_{4}(q^{2})$ by [33, Tab. II] apart from the picky cases identified in Theorem 3.2 only the representatives $u_{9},u_{10},u_{11},u_{12}$ involve one of the simple root elements.
@@
Let $P$ be the standard maximal parabolic subgroup of $G$ of type $A_{n-1}$, that is, the stabiliser of a maximal isotropic subspace in the natural representation, and $Q$ the standard maximal parabolic subgroup of type $B_{n-1}$. We claim that if (up to conjugation) ${\operatorname{Sub}}_{P}(x)=P$ then ${\operatorname{Sub}}_{G}(x)=G$. Indeed, assume ${\operatorname{Sub}}_{P}(x)=P$ and let $\bar{x}$ be the image of $x$ in the standard Levi subgroup $L\cong{\operatorname{GL}}_{n}(q)$ of $P$. By Proposition 3.6, $\bar{x}$ is not regular, so (up to conjugation) its image $y$ in the Levi factor of a standard parabolic subgroup $L_{1}$ of $L$ of type ${\operatorname{GL}}_{n-1}(q)$ is not regular either. Now, $L_{1}$ is a standard Levi subgroup of the standard Levi subgroup $M$ of $Q$ of type ${\operatorname{SO}}_{2n-1}(q)$, and hence by induction ${\operatorname{Sub}}_{M}(y)=M$. (The induction base is given by the case 2n1=52n-1=5 from Proposition 3.7.) But then ${\operatorname{Sub}}_{Q}(x)=Q$, and since $P$ is maximal in $G$ and $Q$ is distinct from $P$, we have ${\operatorname{Sub}}_{G}(x)=G$.
@@
Let $G=F_{4}(q)$ and $x\in G$ be unipotent. Then ${\operatorname{Sub}}_{G}(x)=G$ unless $x$ is regular, or $q$ is odd and $x$ is in class $F_{4}(a_{1})$, with $|{\mathbf{C}}_{G}(x)|=2q^{6}$ and representative $x_{24}$ in [35, Tab. 5 resp.  6], where ${\operatorname{Sub}}_{G}(x_{24})=P_{\{3,4\}}$.
@@
Representatives for the unipotent conjugacy classes of $G$ as well as for a Levi subgroup $L$ of $G$ of type $B_{3}$ were determined in [32, 35]. First assume $p=2$. We claim that only the regular unipotent elements in $L$ have subnormaliser strictly smaller than $L$. Indeed, seven of the class representatives given in [32, Prop. 2.1] do not involve any simple root element, and the other three non-regular ones can be conjugated by a simple reflection to elements $x$ which only involve the 3rd simple root of $L$. Then $x$ has trivial image in the standard Levi subgroup of $L$ of type $B_{2}$, and image involving just one simple root in the standard Levi subgroup of type $A_{2}$, so ${\operatorname{Sub}}_{L}(x)=L$ by Proposition 3.5.
Now we discuss the unipotent class representatives of $G$ from [32, Thm 2.1]. The ones labelled $x_{0}$ through $x_{19}$ do not involve any simple root element, and similarly for $x_{27},x_{28}$. For elements $x\in\{x_{22},\ldots,x_{26}\}$ only one of the short simple roots occurs, so the images of $x$ in standard Levi subgroups of type $A_{2}$ and $B_{3}$ are non-regular and we conclude by Proposition 3.5 and the previous paragraph. Similarly for $x=x_{20},x_{21}$ the image in a Levi of type $C_{3}$ is trivial, and in a Levi of type $B_{3}$ is non-regular. By Lemma 3.8 this leaves us with the elements $x_{29}$ and $x_{30}$, of type $F_{4}(a_{1})$. The given representatives show that these elements are regular in the subsystem subgroup of type $B_{4}$. Conjugating $x=x_{29}$ or $x_{30}$ with the simple reflection not belonging to the $B_{3}$-subsystem, we obtain an element which only involves simple root elements for the 2nd and 3rd simple root. By the first paragraph, the subnormaliser of that element contains the standard parabolic subgroup of type $B_{3}$, but by Proposition 3.6 also the one of type $A_{2}$, hence it is all of $G$.
If $p$ is odd, Proposition 3.7 shows that apart from the regular classes, only the unipotent class of $L$ with representative $z_{8}$ (in the notation of [35, Tab. 3]) has proper subnormaliser. From [35, Tab. 5 and 6] we conclude that again at most the classes of $G$ of type $F_{4}(a_{1})$ with representatives $x_{23},x_{24}$ might have a proper subnormaliser. Conjugating $x=x_{23},x_{24}$ by the simple reflection not in the $B_{3}$-subsystem we obtain elements whose image in $L$ equals $z_{7},z_{8}$ respectively. As observed above, ${\operatorname{Sub}}_{L}(z_{7})=L$ and so ${\operatorname{Sub}}_{G}(x_{23})=G$ by the usual argument. On the other hand, we have ${\operatorname{Sub}}_{L}(z_{8})$ is the standard parabolic subgroup corresponding to the 3rd node of the diagram of $B_{3}$, which is also the 3rd node of the diagram of $F_{4}$. Thus the set of standard parabolic subgroups of $G$ containing $x_{24}$ in their radical has the unique maximal element $P_{\{3,4\}}$, and so ${\operatorname{Sub}}_{G}(x_{24})=P_{\{3,4\}}$ by Lemma 2.14. ∎
@@
Let $G={\operatorname{SU}}_{2m}(q)$ and $x\in G$ be unipotent. Then ${\operatorname{Sub}}_{G}(x)=G$ unless $x$ is regular, or $x$ has Jordan type (2m1,1)(2m-1,1), when ${\operatorname{Sub}}_{G}(x)\sim_{G}P_{m}$.